find the equation of a line that has the given properties. graph each line.
contains (9,3); perpendicular to the line x+3y=-27

Answer by mananth(16946)

1 x + 3 y = -27
Find the slope of this line

3 y = -1 x -27
Divide by 3
y = - 1/ 3 x + -9
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal 3
m1*m2=-1
The slope of the required line will be 3

m= 3 ,point ( 9 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = 27 + b
b= -24
m= 3
The required equation is y=3x-24
m.ananth@hotmail.ca